When pacific inc. Bid for a project with the government the company was offered the following two payment options



a) A payment of $540,000 at the end of 5 years. Which is the scheduled completion time for the project



b) $80,000 paid upfront at the beginning of the project and the balance payment in 5 years.



If the two payments are financially equivalent and the interest rate is 6.00%


Compounded quarterly. Calculate the balance payment offered in ( option b) round to the nearest cent

1
Expert's answer
2022-03-07T03:10:02-0500

"FV=PV(1+\frac{r}{n})^{n×t}"

compounded quarterly n=12,t=5


"=80,000(1+\frac{0.06}{12})^{12×5}"


"=80,000(1.005)^{60}=\$107908"